Circular No. 5930

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PSR 0329+54 R. D. Dagkesamansky, Pushchino Radio Astronomy Observatory; and

Y. P. Shitov, Pulsar Laboratory, Lebedev Physical Institute, report: “Using the pulsar timing technique, T. V. Shabanova has detected a probable planetary system around PSR 0329+54, one of the brightest pulsars. The total data set analyzed by Shabanova extends over 25 yr and consists of Pushchino timing measurements during 1979-Jan. 1994 and some previously published data concerning pulse arrival times (Downs and Reichley 1983, Ap.J. Suppl. 53, 169; Downs and Krause-Polstorff 1986, Ap.J. Suppl. 62, 81). The analysis suggests that PSR 0329+54 is orbited by a planet-like body with an orbital period of ∼ 6140 days. The projected semimajor axis of the pulsar orbit is 17.8 light-ms, and the orbital eccentricity is quite large (0.23). The mass function of the binary system is 1.4 × 10^{−16} M_{⊙}; i.e., assuming that the pulsar mass is 1.4 M_{⊙}, then the companion mass is 2.0 M_{⊕}/sini and has orbital radius ∼ 7.3 AU. Another sinusoidal modulation of the timing residuals with a period ∼ 1110 days and amplitude 1 ms is observed. Assuming the existence of a second planet, the corresponding mass function is 4.5 × 10^{−19} M_{⊙}; i.e., the second planet’s mass is 0.3 M_{⊕}/sini and its orbital radius is ∼ 2.3 AU.”

CYGNUS X-1 B. F. Phlips, Universities Space Research Association and Naval Re-

search Laboratory (NRL); M. D. Leising, Clemson University; K. McNaron-Brown, George Mason University; and J. E. Grove, W. N. Johnson, J. D. Kurfess, and M. S. Strickman, NRL, report for the Compton Observatory OSSE Team: “OSSE is observing Cygnus X-1 between Feb. 1 and 8, based on information from BATSE that the source is in a low state. We measured the integrated 45- to 140-keV flux to be 1.5 and 1.1 (± 0.1) × 10^{−2} photon cm^{−2} s^{−1} for the first two days of the observation, respectively. This is one-fifth of the flux in the low, or γ_{1}, state observed by HEAO-3 (Ling et al. 1987, Ap.J. 321, L117). Optically-thin thermal bremsstrahlung models fit the spectra well and produced temperatures of 56 ± 8 and 62 ± 8 keV for Feb. 1 and 2, respectively. These temperatures are half of the values typically measured by OSSE for Cygnus X-1. We find no evidence for the enhanced 1-MeV emission associated with low states as reported by Ling et al. Observations at other wavelengths are encouraged.”

1994 February 7 Daniel W. E. Green